OpenStudy (woznichak):

can someone please please please help me?? 3x2+5x-3=0 what do i have to take out to it become a quadratic equation

OpenStudy (mathmale):

Nothing. This is a perfectly good quadratic equation as is ... EXCEPT that you must express exponentiation using the " ^ " symbol: 3^x2+5x-3=0

OpenStudy (woznichak):

my teacher gave this to me as a question so i dont know. she said i have to take something out to become something else.

OpenStudy (mathmale):

You'll need to ask your teacher for clarification, then. There is nothing in 3x2+5x-3=0 to "take out." Actually, you need to put in " ^ " to indicate exponentiatin: 3x^2+5x-3=0

OpenStudy (mathmale):

At this point you could try factoring 3x^2+5x-3, which would enable you to "solve for x."